Suppose a firm has a cost curve equal to C 7200+150Q. The firms demand curve is p-550 2Q. (Round all numeric responses to two decimal places,) If regulators set the price equal to the marginal cost, what would be the firms loss? $ -7200 If the regulators set the price equal to the average cost, what would be the price? S 190.00 What would be the deadweight loss in this case of average cost price regulation? $
Answer #1

C = 7200 +150Q

MC = d(C) / dQ = 150

If P = MC = 150

150 = 550 - 2Q

2Q = 400

Q = 200

Economic profit = 200*150 - 7200 - 150*200 = -7200

So firm's loss = 7200 ( or firm's profit = -7200)

- If P = ATC

ATC = C/Q = 7200/Q + 150

P =ATC, so

7200/Q + 150 =550 - 2Q

7200 + 2Q2 =400Q

2Q2 - 400Q + 7200 = 0

solving this we get,

Q = 20 or Q = 180

at Q =20

ATC = 7200 /20 + 150 = $510

and if Q = 180

ATC = 7200/180 + 150 = $190

** Here assume that the regulator set the price = ATC, so the firm only produce Q = 20 ( as at this quantity also they get ATC and at Q = 180 units also they get same price so they stop producing after Q= 20)

P=ATC = $510

DWL ( dead weight loss)

At Q =0, P =550

DWL = (1/2)*( 550- 510)* ( 20) = $400